OpenStudy (18jonea):

Calculate the zeros of the function y=x^2−9x+14 . (0,7) and (0,2) (7,0) and (2,0) (0,14) None of these @calcmathlete

OpenStudy (18jonea):

I think true @Calcmathlete

OpenStudy (anonymous):

Sorry, was helping someone else. To calculate the zeroes, we set y = 0 and factor. \[0 = x^2 - 9x + 14\]So, we need to find two numbers that multiply to 14 but add up to -9. By inspection, -7 and -2 fit the bill. \[0 = x^2 - 9x + 14 \implies 0 = (x - 7)(x - 2)\]Can you get the zeroes from there?

OpenStudy (18jonea):

yes

OpenStudy (anonymous):

To make sure, which of those do you think the answer is?

OpenStudy (18jonea):

B

OpenStudy (anonymous):

Yup! Good job!